OpenStudy (anonymous):

WILL MEDAL AND FAN Given the function f(x) = 0.5(3)^x, what is the value of f^−1(7)? 0.565 1.140 1.771 2.402

OpenStudy (anonymous):

2.402 Im 100% sure

OpenStudy (anonymous):

@julithehero my medal

OpenStudy (anonymous):

Can you tell me how you did it?

OpenStudy (anonymous):

yeah wait

OpenStudy (anonymous):

f(x) = 0.5(3)^x, what is the value of f−1(7)? 7 = 0.5 * (3)^x 7/.5 = (3)^x 14 = 3^x log(14) = x * log(3) x = log(14) / log(3) = 2.402

OpenStudy (anonymous):

no do u get it

OpenStudy (anonymous):

Yeh how do i send you the medal

OpenStudy (anonymous):

just fan me and put best anwser

OpenStudy (anonymous):

best responce

OpenStudy (anonymous):

hoped it helped